If solution A has 10,000 times as many H+ ions as solution B, what is the difference in pH units between the two solutions? Explain.
1
Expert's answer
2019-02-21T10:50:08-0500

Let "a" is concentration of hydrogen ions in solution A, and "b" is concentration of hydrogen ions in solution B.

pH of solution A is log(a), while pH of solution B is log(b).

We know that a/b = 10000

Let's take logarithm of both parts:

log(a/b) = log(10000)

log(a/b) =log(a) - log(b) (it is a known math function), which is the difference in pH we need to identify.

log(a) - log(b) = log(10000)

log(a) - log(b) = 4

So, pH of solution A is 4 units greater than pH of solution B.